OpenStudy (anonymous):

Use the quadratic formula to solve the following quadratic equation: 8x2 + 2x - 3 = 0.

OpenStudy (jdoe0001):

\(\bf \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

here are the answers A. x = -7/8 or 3/8 B. x = -11/8 or 9/8 C. x = -3/4 or 1/2 D. x = -1/2 or 3/4

OpenStudy (jdoe0001):

\(\begin{array}{ccccccclll} &a&&b&&c\\ &8&x^2 &+ 2&x &- 3& = 0 \end{array}\\\quad \\\quad \\ \bf x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (jdoe0001):

plug them in, see what you get

OpenStudy (anonymous):

i dont know what it is. i was taking a test that why i asked

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (jdoe0001):

notice the letters => \(\bf \large \begin{array}{ccccccclll} &\bf a&&\bf b&&\bf c\\ &8&x^2 &+ 2&x &- 3& = 0 \end{array}\)

OpenStudy (jdoe0001):

tis just the coefficients of the quadratic equation

OpenStudy (anonymous):

ohh okay i got it now thanks